a rectangular closet is 2 ft deep, 3ft wide and 8 ft high. what is the length of the longest pole that can fit within the closet?
Answer by stanbon(75887)

a rectangular closet is 2 ft deep, 3ft wide and 8 ft high. what is the length of the longest pole that can fit within the closet?
------
diagonal of the base = sqrt[3^2 + 8^2] = sqrt[73]
-----
longest pole = sqrt[(sqrt(73))^2 + 2^2] = sqrt[73+4] = sqrt(77)
================
Cheers,
Stan H.